1- Billiard ball A with mass 0.81 kg collides head on and elastically with billiard ball B (same mass) at rest. Knowing that Ball A moved at 3.41 m/s before collision, what is the speed of Ball B after the collision?

2- A 4.78 kg cube sits on a frictionless surface. A 11.99 gram bullet traveling at 388.07 m/s is shot into the brick and they move together at a certain speed. What is the speed in m/s?

3-  What is the SI unit for momentum?
Answer #1

1)

In elastic collision if object 2 is at rest.

v2f = 2*m1*v1i / (m1+m2)

v1f = (m1-m2)*v1i/(m1+m2)

when m1=m2

v2f = v1i

   = 3.41 m/s

Answer: 3.41 m/s

2)

use conservation of momentum

initial momentum = final momentum

m1*vi = (m1+m2)*vf

11.99*10^-3 kg * 388.07 m/s = (11.99*10^-3 + 4.78)*vf

vf = 0.971 m/s

Answer: 0.971 m/s

3)

SI unit of momentum is

Kgm/s